# CAUSAL INFERENCE ON PROCESS GRAPHS PART II CAUSAL STRUCTURE AND EFFECT IDENTIFICATION

### v1

@@ Appendix B Proof of the spectral Gessel-Viennot Lemma
$\mathfrak{m}_{1,j}$ $=\sum_{\Pi\in\mathbf{N}_{0}(X_{1},Y_{j})}\mathrm{sgn}(\Pi)\prod_{\pi\in\Pi}h^{\mathrm{P}(\pi)}$
@@
$\mathfrak{h}^{\infty}_{x_{1},y_{j}}$ $=\sum_{\pi\in\mathbf{P}_{0}(x_{1},y_{j})}h^{\mathrm{P}(\pi)}.$
@@
$\det([H^{\infty}]_{X,Y})$ $=\sum_{j=1}^{l}\sum_{\pi\in\mathbf{P}_{0}(x_{1},y_{j})}\sum_{\Pi\in\mathbf{N}_{0}(X_{1},Y_{j})}\mathrm{sgn}(\Pi\cup\pi)h^{\mathrm{P}(\pi)}H^{\mathrm{P}(\Pi)}$ (13)
@@
Then $\Pi^{\prime}=\Pi\setminus\{\pi^{\prime}\}\cup\{\tilde{\pi}\}\in\mathbf{P}_{0}(X_{1},Y_{k})$ is a system of non intersecting directed paths from $X_{1}$ to $Y_{k}$. That the system is non intersecting follows from the choice of $v$ and $\pi^{\prime}$. Furthermore, $\hat{\pi}$ is a directed path from $x_{1}$ to $y_{j}$. It follows that
$h^{\mathrm{P}(\pi)}H^{\mathrm{P}(\Pi)}$ $=h^{\mathrm{P}(\hat{\pi})}H^{\mathrm{P}(\Pi^{\prime})}$
@@
It remains to show that $\det([I-H]_{X,Y}^{-1})$ is non-zero if and only if there there is at least one system of non-intersecting paths from $X$ to $Y$. First, if there is no such system then it follows by the identity we just showed that the determinant must be zero. On the other hand, if $\mathbf{P}_{0}(X,Y)$ is non-empty, then it must be finite (because $V$ is finite). So suppose $\mathbf{P}_{0}(X,Y)=\{\Pi_{1},\dots,\Pi_{l}\}$. We now wish to show that the rational function
$H^{\mathrm{P}(\Pi_{1})}+\cdots+H^{\mathrm{P}(\Pi_{l})}\in\mathbb{R}(\Phi,z)\setminus\{0\},$ (14)
i.e. that this is a non-zero rational function in the indeterminates $\Phi$ and $z$. In order to show this claim we introduce a couple of polynomials. Let $\pi$ be a directed path, then
@@
of which the first is the numerator and the second is denominator of the path function $h^{(\pi)}$. These are polynomials in $\mathbb{R}[\Phi,z]$ and note that the coefficient of the degree zero monomial of $q^{(\pi)}$ is equal to one. Now suppose that $\pi\in\mathbf{P}_{0}(x,y)$ is a directed path that visits every vertex at most once and $\Pi\in\mathbf{P}_{0}(X,Y)$ a system of directed paths, then we introduce the following polynomials
$p^{\mathrm{P}(\pi)}$ $=p^{(\pi)}\prod_{\zeta\in\mathcal{C}_{\sim}(\pi)}q^{(\zeta)}$ $q^{\mathrm{P}(\pi)}$ $=q^{(\pi)}\prod_{\zeta\in\mathcal{C}_{\sim}(\pi)}q^{(\zeta)}-p^{(\zeta)}$
$p^{\mathrm{P}(\Pi)}$ $=\prod_{\pi\in\Pi}p^{\mathrm{P}(\pi)}$ $q^{\mathrm{P}(\Pi)}$ $=\prod_{\pi\in\Pi}q^{\mathrm{P}(\pi)}.$
@@
$p^{\mathrm{P}(\Pi_{1})}Q_{1}+\dots+p^{\mathrm{P}(\Pi_{l})}Q_{l}\in\mathbb{R}[\Phi,z],$ $Q_{i}$ $=\prod_{j\neq i}^{l}q^{\mathrm{P}(\Pi_{j})}\in\mathbb{R}[\Phi,z]$
@@
Note that the coefficient of $x$ in $p^{\mathrm{P}(\Pi_{1})}Q_{1}$ is equal to one, which can be seen from the fact that the coefficient of the monomial with degree zero is equal to one in every $Q_{i}$. It remains to show that the coefficient of $x$ in $p^{\mathrm{P}(\Pi_{j})}Q_{j}$ is zero for all other $j$. Let us therefore consider the sets of links $D_{i}=\{v\to w\mid v\to w\in\pi,\pi\in\Pi_{i}\}$. Note that for any two distinct $i,j$ the assumption $\Pi_{i}\neq\Pi_{j}$ are systems of non-intersecting paths implies that also the corresponding sets of edges are different, i.e. $D_{i}\neq D_{j}$. This implies that the monomial $x$ has a non-zero coefficient only in the polynomial $p^{\mathrm{P}(\Pi_{1})}Q_{1}$. This allows us to conclude that the rational function (14) is indeed non-zero. ∎

### v2

@@ Appendix B Proof of the spectral Gessel-Viennot Lemma
$\mathfrak{m}_{1,j}$ $=\sum_{\Pi\in\mathbf{N}_{0}(X_{1},Y_{j})}\mathrm{sgn}(\Pi)\prod_{\pi\in\Pi}h^{(\pi)}$
@@
$\mathfrak{h}^{\infty}_{x_{1},y_{j}}$ $=\sum_{\pi\in\mathbf{P}_{0}(x_{1},y_{j})}h^{(\pi)}.$
@@
$\det([H^{\infty}]_{X,Y})$ $=\sum_{j=1}^{l}\sum_{\pi\in\mathbf{P}_{0}(x_{1},y_{j})}\sum_{\Pi\in\mathbf{N}_{0}(X_{1},Y_{j})}\mathrm{sgn}(\Pi\cup\pi)h^{(\pi)}H^{(\Pi)}$ (16)
@@
Then $\Pi^{\prime}=\Pi\setminus\{\pi^{\prime}\}\cup\{\tilde{\pi}\}\in\mathbf{P}(X_{1},Y_{k})$ is a system of non intersecting directed paths from $X_{1}$ to $Y_{k}$. That the system is non intersecting follows from the choice of $v$ and $\pi^{\prime}$. Furthermore, $\hat{\pi}$ is a directed path from $x_{1}$ to $y_{j}$. It follows that
$h^{(\pi)}H^{(\Pi)}$ $=h^{(\hat{\pi})}H^{(\Pi^{\prime})}$
@@
It remains to show that $\det([I-H]_{X,Y}^{-1})$ is non-zero if and only if there there is at least one system of non-intersecting paths from $X$ to $Y$. First, if there is no such system then it follows by the identity we just showed that the determinant must be zero. So suppose $\mathbf{P}(X,Y)=\{\Pi_{1},\dots,\Pi_{l}\}$. We now wish to show that the rational function
$H^{(\Pi_{1})}+\cdots+H^{(\Pi_{l})}\in\mathbb{R}(\Phi,z)\setminus\{0\},$ (17)
i.e. that this is a non-zero rational function in the indeterminates $\Phi$ and $z$. In order to show this claim we introduce a few polynomials. Let $\pi$ be a directed path, then
@@
of which the first is the numerator and the second is denominator of the path function $h^{(\pi)}$. These are polynomials in $\mathbb{R}[\Phi,z]$ and note that the coefficient of the degree zero monomial of $q^{(\pi)}$ is equal to one. Now suppose that $\pi\in\mathbf{P}_{0}(x,y)$ is a directed path that visits every vertex at most once and $\Pi\in\mathbf{P}(X,Y)$ a system of directed paths, then we introduce the following polynomials
$p^{(\Pi)}$ $=\prod_{\pi\in\Pi}p^{(\pi)}$ $q^{(\Pi)}$ $=\prod_{\pi\in\Pi}q^{(\pi)}.$
@@
$p^{(\Pi_{1})}Q_{1}+\dots+p^{(\Pi_{l})}Q_{l}\in\mathbb{R}[\Phi,z],$ $Q_{i}$ $=\prod_{j\neq i}^{l}q^{(\Pi_{j})}\in\mathbb{R}[\Phi,z]$
@@
Note that the coefficient of $x$ in $p^{(\Pi_{1})}Q_{1}$ is equal to one, which can be seen from the fact that the coefficient of the monomial with degree zero is equal to one in every $Q_{i}$. It remains to show that the coefficient of $x$ in $p^{(\Pi_{j})}Q_{j}$ is zero for all other $j$. Let us therefore consider the sets of links $D_{i}=\{v\to w\mid v\to w\in\pi,\pi\in\Pi_{i}\}$. Note that for any two distinct $i,j$ the assumption $\Pi_{i}\neq\Pi_{j}$ are systems of non-intersecting paths implies that also the corresponding sets of edges are different, i.e. $D_{i}\neq D_{j}$. This implies that the monomial $x$ has a non-zero coefficient only in the polynomial $p^{(\Pi_{1})}Q_{1}$. This allows us to conclude that the rational function (17) is indeed non-zero. ∎
Appendix C Auxiliary Lemma for proving the LF-HTC identifiability result
Suppose $\mathcal{G}$ is a time series graph with exogenous latent structure. In the main part of the paper we introduced $\mathbb{R}^{\mathcal{D}}_{\mathrm{stb}}$ as the space of SVAR parameters that are compatible with $\mathcal{G}$ and give rise to process which can be represented as an SEP in the frequency domain. The following Lemma is used in the main part of the paper to prove the spectral version of the latent factor half-trek identifiability result. The Lemma is closely related to Lemma 1 in the supplementrary material of [16].
Lemma 6.
Suppose $X,Y\subset O$ are disjoint subsets and let $T^{\prime}=(\tau_{1},\dots,\tau_{n})\in\mathbf{T}(X,Y)$ be a system of latent factor half-treks without sided intersection. Then there exists a sub-system $T=(\tau_{1},\dots,\tau_{n})\in\mathbf{T}(X,Y)$ of latent factor half-treks together with an ordering of the elements in $X$ and $Y$ and $T$ such that:
  1. 1.
    $\tau_{i}\in\mathbf{T}(x_{i},y_{i})$ a latent factor half-trek such that $\tau_{i}$ goes through $x_{j}$ only if $i\geq j$ and $x_{i}$ appears on $\tau_{i}$ exactly once
  2. 2.
    the associated subgraph $G(T)$ is acyclic.
For such $T$ and associated parameter $\Phi\in\mathbb{R}^{\mathcal{D}(T)}_{\mathrm{stb}}$ it holds that
$\det([\mathcal{S}_{\mathcal{G}(T)}(\Phi,\Omega)]_{X,Y})=\mathcal{S}^{(T)}(\Phi,\Omega)$ (18)
Proof.
Using the same procedure as in the proof of Lemma 1 in the supplementary material of [16], one can reduce the system $T^{\prime}$ to a system of latent factor half-treks $T\in\mathbf{T}(X,Y)$ with property 2.) and furthermore reorder the elements in $X$, $Y$ and $T$ so that 1.) holds.
To show the equation (18) we need to show that $\mathbf{T}_{G(T)}(X,Y)=\{T\}$. Let $T^{\prime}=(\tau_{1}^{\prime},\dots,\tau_{n}^{\prime})$ be a trek system without sided intersection from $X$ to $Y$ such that $\tau^{\prime}_{i}\in\mathbf{T}(x_{i},y_{\alpha(i)})$, where $\alpha$ is a permutation on $\{1,\dots,n\}$. We need to show that $T^{\prime}=T$.
Each $\tau_{k}$ is a latent factor half-trek, so it has one of the following two forms
$x_{k}\leftarrow l_{k}\to z_{1}^{k}\to\cdots z^{k}_{I_{k}}=y_{k}$ $x_{k}\to z_{1}^{k}\to\cdots z^{k}_{I_{k}}=y_{k},$
where $l_{k}\in L$ and $z_{j}^{k}\in O$. Each $x\in X$ has at most one outgoing link on $G(T)$, this follows from the fact that $T$ has no side intersection and $G(T)$ is acyclic.
We start by showing that $\tau_{n}^{\prime}=\tau_{n}$. To construct a trek $\tau^{\prime}_{n}$ starting at $x_{n}$ we need an edge in $D(T)$ that either points to $x_{n}$ or emerges from $x_{n}$. By assumption $x_{n}$ appears only on $\tau_{n}$ and on $\tau_{n}$ the vertex $x_{n}$ appears exactly once. So there is exactly one edge involving $x_{n}$, which means that the first link on $\tau_{n}^{\prime}$ must be the first link on $\tau_{n}$.
If the first link is $x_{n}\leftarrow l_{n}$, then the second link on $\tau_{n}^{\prime}$ cannot be $l_{n}\to x_{n}$ again, because $X$ and $Y$ are disjoint and there is no path from $x_{n}$ to any $y\in Y$ on $G(T)$. Since $l_{n}$ has exactly two emerging links on $G(T)$, the second link on $\tau_{n}^{\prime}$ must be $l_{n}\to z_{1}$. In particular, if $z_{1}\in Y$, then $z_{1}=y_{n}$. If $z_{1}\neq y_{n}$, then there can only be one edge coming out of $z_{1}$ on $G(T)$, since $T$ has no sided intersection. So the next edge on $\tau_{n}^{\prime}$ is $z_{1}\to z_{2}$. This reasoning can be repeated until $y_{n}$ is reached. The case where $\tau_{n}=x_{n}\to z_{1}\to\cdots$ works analogously. So, we conclude that $\tau_{n}^{\prime}=\tau_{n}$.
Now suppose we have shown that $\tau_{k}=\tau_{k}^{\prime}$ for all $m<k\leq n$. We now want to show that $\tau_{m}^{\prime}=\tau_{m}$. By the assumption on $T$, the node $x_{m}$ can only appear on $\tau_{j}$ with $j\geq m$. To construct a trek starting at $x_{m}$ we need either a link pointing to $x_{m}$ or a link coming from $x_{m}$.
Case 1: If $x_{m}$ only appears on $\tau_{m}$, then we can repeat the same argument as above to show that $\tau_{m}^{\prime}=\tau_{m}$.
Case 2: The node $x_{m}$ also appears on some $\tau_{l}=\tau_{l}^{\prime}$ with $l>m$, i.e. $x_{m}=z^{l}_{j}$ for a $j\in\{2,\dots,I_{l}-1\}$, where $l$ and $j$ are unique, since $T$ is without sided intersection and $G(T)$ is acyclic. In this case there can be only one edge on $G(T)$ starting from $x_{m}$, i.e. $x_{m}=z^{l}_{j}\to z^{l}_{j+1}$. By assumption, this edge already occurs at $\tau_{l}^{\prime}=\tau_{l}$. So the first edge on the path $\tau_{l}^{\prime}$ must point to $x_{m}$ (otherwise $\tau_{m}^{\prime}$ would have a sided intersection with $\tau_{l}^{\prime}$). There are now exactly two links we can choose to start $\tau^{\prime}_{m}$, i.e. $x_{m}\leftarrow l_{m}$ or $z\to z_{j}^{l}$, where $z$ is either $z_{j-1}^{l}$ or $l_{l}$ or $x_{l}$. We now need to show that the first link on $\tau^{\prime}_{m}$ is indeed $x_{m}\leftarrow l_{m}$. Once we have shown that, we can proceed as above to prove that $\tau_{m}^{\prime}=\tau_{m}$, since $\tau_{m}^{\prime}$ cannot be of the form $x_{m}\leftarrow l_{m}\to x_{m}$ (then $\tau_{m}^{\prime}$ would intersect with $\tau_{l}^{\prime}$).
Suppose $j>1$ and also that there are some $1\leq i<j$ such that $z_{i}^{l}=x_{q}\in X$. Otherwise, a trek $\tau^{\prime}_{m}$ where the first link was $x_{m}\leftarrow z$ could not reach a point in $Y$, contradicting $T^{\prime}\in\mathbf{T}(X,Y)$.
So let $i^{\ast}=\min\{i\mid i<j,z^{l}_{i}\in X\}$ and $z_{i^{\ast}}^{l}=x_{q}$, so the first link on $\tau_{q}$ is $x_{q}\leftarrow l_{q}$. In this case $\tau_{m}^{\prime}=(x_{m}\leftarrow\cdots\leftarrow x_{q}\leftarrow\cdots)$. Given that $\tau_{m}^{\prime}$ ends in $Y$, there are two options for $\tau^{\prime}_{m}$, namely
$x_{m}\leftarrow\cdots x_{q}\leftarrow l_{q}\to x_{q}$ (19)
$x_{m}\leftarrow\cdots x_{q}\leftarrow l_{q}\to z_{1}^{q}$ (20)
The trek $\tau_{m}^{\prime}$ cannot be of the form (20), since this would result in an intersection with $\tau_{l}$. Nor can $\tau_{m}^{\prime}$ be of the form (19), because that would mean that either $\tau_{m}^{\prime}$ or $\tau_{l}^{\prime}$ would have a sided intersection with the trek $\tau_{q}^{\prime}$ (i.e. the trek in $T^{\prime}$ starting at $x_{q}$). Since this argument can be repeated for every $i<j$ such that $z_{i}^{l}\in X$, we conclude that the first link on $\tau_{m}^{\prime}$ is indeed $x_{m}\leftarrow l_{m}$. As mentioned above, this implies that $\tau_{m}^{\prime}=\tau_{m}$.
This concludes the proof for $\mathbf{T}(X,Y)=\{T\}$ implying equation (18). ∎